OpenStudy (anonymous):

Find the value of square root(72 + square root(72 + square root(72 + (...))) How do I do this?

OpenStudy (anonymous):

If y=sqrt(72+sqrt(72+........) then y=sqrt(72+y). As starting from the second square root the series is the same as y again. y^2=72+y y^2-y-72=0. y-9)(y+8)=0. that means y=9.

OpenStudy (anonymous):

Thank you